1. Xun, Yue and Zhuo have interests in XYZ Partnership. Partnership income for the year is $350,000. The partnership agreement specifies that Xun is to receive an annual salary of $200,000, Yue is to receive an annual salary of $50,000, and Zhuo is to receive an annual salary of $120,000. Any remaining income or loss is to be divided between the three partners in a 2:1:1 ratio. Salaries are to be fully implemented. Partnership income allocated to Yue is:

a.            $40,000

b.            $55,000

c.             $50,000

d.            $45,000

2.       Xun, Yue and Zhuo have interests in XYZ Partnership. Partnership income for the year is $296,000. The partnership agreement specifies that Xun is to receive an annual salary of $200,000, Yue is to receive an annual salary of $50,000, and Zhuo is to receive an annual salary of $120,000. Any remaining income or loss is to be divided between the three partners in a 2:1:1 ratio. Salaries are to be proportionately implemented. Partnership income allocated to Xun is:

a.            $172,000

b.            $163,000

c.             $160,000

d.            $200,000
Answer #1

1.)

Partnership Income for the year = $3,50,000

Salaries to be distributed fully = $3,70,000

(2,00,000+50,000+1,20,000)

Loss to be divided = ​​​​​​350000 -370000 = $20,000

Yue share in the loss (1/4) = ($5,000)

Yue salary Income = $50,000

Net Income received by Yue from Partnership firm = 50000-5000 = $ 45,000

Answer (d) $45,000

2)Since in the second case salary is proportional implemented, so salary shall be limited to available profits.

so, maximum salary that can be distributed is $2,96,000 , this will be divided among partner in the ration of their salary(fixed) which is 20:12:5

So, Xun share in the income will be $2,96,000*20/37 = $1,60,000

Answer (c) i.e. $1,60,000